### v1

# ODGEN: Domain-specific Object Detection Data Generation with Diffusion Models

Abstract
Modern diffusion-based image generative models have made significant progress and become promising to enrich training data for the object detection task. However, the generation quality and the controllability for complex scenes containing multi-class objects and dense objects with occlusions remain limited. This paper presents ODGEN, a novel method to generate high-quality images conditioned on bounding boxes, thereby facilitating data synthesis for object detection. Given a domain-specific object detection dataset, we first fine-tune a pre-trained diffusion model on both cropped foreground objects and entire images to fit target distributions. Then we propose to control the diffusion model using synthesized visual prompts with spatial constraints and object-wise textual descriptions. ODGEN exhibits robustness in handling complex scenes and specific domains. Further, we design a dataset synthesis pipeline to evaluate ODGEN on 7 domain-specific benchmarks to demonstrate its effectiveness. Adding training data generated by ODGEN improves up to 25.3% mAP@.50:.95 with object detectors like YOLOv5 and YOLOv7, outperforming prior controllable generative methods. In addition, we design an evaluation protocol based on COCO-2014 to validate ODGEN in general domains and observe an advantage up to 5.6% in mAP@.50:.95 against existing methods.
Figure 1: The proposed ODGEN enables controllable image generation from bounding boxes and text prompts. It can generate high-quality data for complex scenes, encompassing multiple categories, dense objects, and occlusions, which can be used to enrich the training data for object detection.
1 Introduction
Object detection is one of the most widely used computer vision tasks in real-world applications. However, data scarcity poses a significant challenge to the performance of state-of-the-art models like YOLOv5 [23] and YOLOv7 [52]. With the progress of generative diffusion models, such as Stable Diffision [42], and DALL-E [12], which can generate high-quality images from text prompts, recent works have started to explore the potential of synthesizing images for perceptual model training. Furthermore, methods like ControlNet [60], ReCo [59], and GLIGEN [31] provide various ways to control the contents of synthetic images. Concretely, extra visual or textual conditions are introduced as guidance for diffusion models to generate objects under certain spatial constraints. Therefore, it becomes feasible to generate images together with instance-level or pixel-level annotations.
Nevertheless, it remains challenging to generate an effective supplementary training set for real-world object detection applications. Firstly, large-scale pre-trained diffusion models are usually trained on web crawl datasets such as LAION [46], whose distributions may be quite distinct from specialist domains. The domain gap could undermine the fidelity of generated images, especially the detailed textures and styles of the objects, e.g., the tumor in medical images or the avatar in a video game. Secondly, the text prompts for object detection scenes may contain multiple categories of subjects. It could cause the "concept bleeding" [38] problem for modern diffusion models, which is defined as the unintended merging or overlapping of distinct visual elements in an image. As a result, the synthetic images can be inconsistent with pseudo labels used as generative conditions. Thirdly, overlapping objects, which are common in object detection datasets, are likely to be merged as one single object by existing controllable generative methods. Lastly, some objects may be neglected by the diffusion model so that no foreground is generated at the conditioned location. All of the limitations can hinder the downstream object detection performance.
To address these challenges, we propose ODGEN, a novel method to generate synthetic images with a fine-tuned diffusion model and object-wise conditioning modules that control the categories and locations of generated objects. We first fine-tune the diffusion model on a given domain-specific dataset, with not only entire images but also cropped foreground patches, to improve the synthesis quality of both background scene and foreground objects. Secondly, we design a new text prompt embedding process. We propose to encode the class name of each object separately by the frozen CLIP [39] text encoder, to avoid mutual interference between different concepts. Then we stack the embeddings and encode them with a newly introduced trainable text embedding encoder. Thirdly, we propose to use synthetic patches of foreground objects as the visual condition. Each patch is resized and pasted on an empty canvas according to bounding box annotations, which deliver both conceptual and spatial information without interference from overlap between objects. In addition, we train foreground/background discriminators to check whether the region of each pseudo label contains a synthetic object. If no object can be found, the pseudo label will be filtered.
We evaluate the effectiveness of our ODGEN in specific domains on 7 subsets of the Roboflow-100 benchmark [6]. Extensive experimental results show that adding our synthetic data improves up to 25.3% mAP@.50:.95 on YOLO detectors, outperforming prior controllable generative methods. Furthermore, we validate ODGEN in general domains with an evaluation protocol designed based on COCO-2014 [35] and gain an advantage up to 5.6% in mAP@.50:.95 than prior methods.
The main contributions of this work can be summarized as follows:
  • We propose to fine-tune pre-trained diffusion models with both cropped foreground patches and entire images to generate high-quality domain-specific target objects and background scenes.
  • We design a novel strategy to control diffusion models with object-wise text prompts and synthetic visual conditions, improving their capability of generating and controlling complex scenes.
  • We conduct extensive experiments to demonstrate that our synthetic data effectively improves the performance of object detectors. Our method outperforms prior works on fidelity and trainability in both specific and general domains.
2 Related Works
Diffusion models [47, 19, 48, 7, 36, 27] have made great progress in recent years and become mainstream for image generation, showing higher fidelity and training stability than prior generative models like GANs [15, 4, 25, 26, 24] and VAEs [28, 41, 50]. Latent diffusion models (LDMs) [42] perform the denoising process in the compressed latent space and achieve flexible generators conditioned on inputs like text and semantic maps. LDMs significantly improve computational efficiency and enable training on internet-scale datasets [46]. Modern large-scale text-to-image diffusion models including Imagen [44], DALL·E [40], eDiff-I [2], and Stable Diffusion [42, 38] have demonstrated unparalleled capabilities to produce diverse samples with high fidelity given unfettered text prompts.
Layout-to-image generation synthesizes images conditioned on graphical inputs of layouts. Pioneering works [33, 20] based on GANs [15] and transformers [51] successfully generate images consistent with given layouts. LayoutDiffusion [64] fuses layout and category embeddings and builds object-aware cross-attention for local conditioning with traditional diffusion models. LayoutDiffuse [22] employs layout attention modules for bounding boxes based on LDMs. MultiDiffusion [3] and BoxDiff [56] develop training-free frameworks to produce samples with spatial constraints. GLIGEN [31] inserts trainable gated self-attention layers to pre-trained LDMs to fit specific tasks and is hard to generalize to scenes uncommon for pre-trained models. ReCo [59] and GeoDiffusion [5] extend LDMs with position tokens added to text prompts and fine-tunes both text encoders and diffusion models to realize region-aware controls. They need abundant data to build the capability of encoding the layout information in text prompts. ControlNet [60] reuses the encoders of LDMs as backbones to learn diverse controls. However, it still struggles to deal with some complex cases. MIGC [65] decomposes multi-instance synthesis to single-instance subtasks utilizing additional attention modules. InstanceDiffusion [53] adds precise instance-level controls, including boxes, masks, and scribbles for text-to-image generation. Given annotations of bounding boxes, this paper introduces a novel approach applicable to both specific and general domains to synthesize high-fidelity complex scenes consistent with annotations for the object detection task.
Dataset synthesis for training object detection models. Copy-paste [9] is a simple but effective data augmentation method for detectors. Simple Copy-Paste [13] achieves improvements with a simple random placement strategy for objects. Following works [12, 63] generate foreground objects and then paste them on background images. However, generated images by these methods usually have obvious artifacts on the boundary of pasted regions. DatasetGAN [62] takes an early step to generate labeled images automatically. Diffusion models have been used to synthesize training data and benefit downstream tasks including object detection [11, 5, 61, 10], image classification [1, 8, 17, 32, 45, 49], and semantic segmentation [54, 21, 14, 29, 37, 55, 58, 34, 57]. Modern approaches for image-label pairs generation can be roughly divided into two categories. One group of works [14, 54, 11, 29, 34, 57, 61] first generate images and then apply a pre-trained perception model to generate pseudo labels on the synthetic images. The other group [5, 37, 21] uses the same strategy as our approach to synthesize images under the guidance of annotations as input conditions. The second group also overlaps with layout-to-image approaches [31, 60, 59, 65, 53, 22, 64, 30]. Our approach should be assigned to the second group and is designed to address challenging cases like multi-class objects, occlusions between objects, and specific domains.
Figure 2: ODGEN training pipeline: (a) A pre-trained diffusion model is fine-tuned on a detection dataset with both entire images and cropped foreground patches. (b) A text list is built based on class labels. The fine-tuned diffusion model in stage (a) is used to generate a synthetic object image for each text. Generated object images are resized and pasted on empty canvases per box positions, constituting an image list. (c) The image list is concatenated in the channel dimension and encoded as conditions for ControlNet. The text list is encoded by the CLIP text encoder, stacked, and encoded again by the text embedding encoder as inputs for ControlNet.
3 Method
This section presents ODGEN, a novel approach to generate high-quality images conditioned on bounding box labels for specific domains. Firstly, we propose a new method to fine-tune the diffusion model in Sec. 3.1. Secondly, we design an object-wise conditioning method in Sec. 3.2. Finally, we introduce a generation pipeline to build synthetic datasets for detector training enhancement in Sec. 3.3.
3.1 Domain-specific Diffusion Model Fine-tuning
Modern Stable Diffusion [42] models are trained on the LAION-5B [46] dataset. Therefore, the textual prompts and generated images usually follow similar distributions to the web crawl data. However, real-world object detection datasets could come from very specific domains. We fine-tune the UNet of the Stable Diffusion to fit the distribution of an arbitrary object detection dataset. For training images, we use not only the entire images from the dataset but also the crops of foreground objects. In particular, we crop foreground objects and resize them to $512\times 512$. In terms of the text input, as shown in Fig. 2 (a), we use templated condition "a <scene>" named $c_{s}$ for entire images and "a <classname>" named $c_{o}$ for cropped objects. If the names are terminologies, following the subject-driven approach in DreamBooth [43], we can employ identifiers like "[V]" to represent the target objects and scenes, improving the robustness to textual ambiguity under the domain-specific context. $x_{s}^{t}$ and $x_{o}^{t}$ represent the input of scenes $x_{s}$ and objects $x_{t}$ added to noises at time step $t$. $\epsilon_{s}$ and $\epsilon_{o}$ represent the added noises following normal Gaussian distributions. The pre-trained Stable Diffusion parameterized by $\theta$ is fine-tuned with the reconstruction loss as:
$\mathcal{L}_{rec}=$ $\mathbb{E}_{x_{o},t,\epsilon_{o}\sim\mathcal{N}(0,1)}\left[||\epsilon_{o}-\epsilon_{\theta}(x_{o}^{t},t,\tau(c_{o}))||^{2}\right]$ (1)
$+\lambda\mathbb{E}_{x_{s},t,\epsilon_{s}\sim\mathcal{N}(0,1)}\left[||\epsilon_{s}-\epsilon_{\theta}(x_{s}^{t},t,\tau(c_{s}))||^{2}\right]$
where $\lambda$ controls the relative weight for the reconstruction loss of scene images. $\tau$ is the frozen CLIP text encoder. Our approach guides the fine-tuned model to capture more details of foreground objects and maintain its capability of synthesizing the complete scenes.
3.2 Object-wise Conditioning
ControlNet [60] can perform controllable synthesis with visual conditions. However, the control can be challenging with an increasing object number and category number due to the "concept bleeding" phenomenon [38]. Stronger conditions are needed to ensure high-fidelity generation. To this end, we propose a novel object-wise conditioning strategy with ControlNet.
Text list encoding. As shown in Fig. 2 (b) and (c), given a set of object class names and their bounding boxes, we first build a text list consisting of each object with the fixed template "a <classname>". Then the text list is padded with empty texts to a fixed length $N$ and converted to a list of embeddings by a pre-trained CLIP text tokenizer and encoder. The embeddings are stacked and encoded by a 4-layer convolutional text embedding encoder, and used as the textual condition for ControlNet. The text encoding of native ControlNet compresses the information in global text prompts altogether, resulting in mutual interference between distinct categories. To alleviate this "concept bleeding" [38] problem of multiple categories, our two-step text encoding enables the ControlNet to capture the information of each object with separate encoding.
Image list encoding. As shown in Fig. 2 (b) and (c), for each item in the text list, we use the fine-tuned diffusion model to generate images for each foreground object. We then resize each generated image and paste it on an empty canvas based on the corresponding bounding box. The set of pasted images is denoted as an image list, which contains both conceptual and spatial information of the objects. The image list is concatenated and zero-padded to $N$ in the channel dimension and then encoded to the size of latent space by an image encoder. Applying an image list rather than pasting all objects on a single image can effectively avoid the influence of object occlusions.
Optimization. With a pair of image and object detection annotation, we generate the text list $c_{tl}$ and image list $c_{il}$ as introduced in this section. The input global text prompt $c_{t}$ of the diffusion model is composed of the object class names and a scene name, which is usually related to the dataset name and fixed for each dataset. The ControlNet, along with the integrated encoders, are trained with the reconstruction loss. In addition, the weights of foreground regions are enhanced to produce more realistic foreground objects in complex scenes. The overall loss function is formulated as:
$\begin{gathered}\mathcal{L}_{recon}=\mathbb{E}_{x,t,c_{t},c_{tl},c_{il},\epsilon\sim\mathcal{N}(0,1)}\left[||\epsilon-\epsilon_{\theta}(x_{t},t,c_{t},c_{tl},c_{il})||^{2}\right]\\ \mathcal{L}_{control}=\mathcal{L}_{recon}+\gamma\mathcal{L}_{recon}\odot\mathcal{M}\end{gathered}$ (2)
where $\mathcal{M}$ represents a binary mask with $1$ on foreground pixels 00 on background pixels. $\odot$ represents element-wise multiplication, and $\gamma$ controls the foreground re-weighting.
Figure 3: Pipeline for object detection dataset synthesis. Yellow block: estimate Gaussian distributions for the bounding box number, area, aspect ratio, and location based on the training set. Blue block: sample pseudo labels from the Gaussian distributions and generate conditions including text and image lists to synthesize novel images. Pink block: train a classifier with foreground and background patches randomly cropped from the training set and use it to filter pseudo labels that failed to be synthesized. Finally, the filtered labels and synthetic images compose datasets.
3.3 Dataset Synthesis Pipeline for Object Detection
Our dataset synthesis pipeline is summarized in Fig. 3. We generate random bounding boxes and classes as pseudo labels based on the distributions of the training set. Then, the pseudo labels are converted to triplets: <image list, text list, global text prompt>. ODGEN uses the triplets as inputs to synthesize images with the fine-tuned diffusion model. In addition, we filter out the pseudo labels in which areas the foreground objects fail to be generated.
Object distribution estimation. To simulate the training set distribution, we compute the mean and variance of bounding box attributes and build normal distributions. In particular, given a dataset with $K$ categories, we calculate the class-wise average occurrence number per image $\boldsymbol{\mu}=(\mu_{1},\mu_{2},\dots,\mu_{K})$ and the covariance matrix $\boldsymbol{\Sigma}$ to build a multi-variable joint normal distribution $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. In addition, for each category $k$, we build normal distributions $\mathcal{N}(\mu_{x},\sigma^{2}_{x})$ and $\mathcal{N}(\mu_{y},\sigma^{2}_{y})$ for the top-left coordinate of the bounding boxes, $\mathcal{N}(\mu_{area},\sigma^{2}_{area})$ for box areas, and distributions $\mathcal{N}(\mu_{ratio},\sigma^{2}_{ratio})$ for aspect ratios.
Image synthesis. We first sample the object number per category from the $K$-dimensional joint normal distributions. Then for each object, we sample its bounding box location and size from the corresponding estimated normal distributions. The sampled values are used to generate text lists, image lists, and global text prompts. Finally, we use ODGEN to generate images with the triplets.
Corrupted label filtering. There is a certain possibility that some objects fail to be generated in synthetic images. While utilizing image lists alleviates the concern to some extent, the diffusion model may still neglect some objects in complex scenes containing dense or overlapping objects. As a result, the synthesized pixels could not match the pseudo labels and will undermine downstream training performance. We fine-tune a ResNet50 [16] with foreground and background patches cropped from the training set to classify whether the patch contains an object. The discriminator checks whether objects are successfully synthesized in the regions of pseudo labels, rejects nonexistent ones, and further improves the accuracy of synthetic datasets.
Figure 4: Comparison between ODGEN and other methods under the same condition shown in the first column. ODGEN can be generalized to specific domains and enables accurate layout control.
Table 1: FID ($\downarrow$) scores computed over 5000 images synthesized by each approach on RF7 datasets. ODGEN achieves better results than the other on all 7 domain-specific datasets.
Datasets ReCo GLIGEN ControlNet GeoDiffusion ODGEN
Apex Game 88.69 125.27 97.32 120.61 58.21
Robomaster 70.12 167.44 134.92 76.81 57.37
MRI Image 202.36 270.52 212.45 341.74 93.82
Cotton 108.55 89.85 196.87 203.02 85.17
Road Traffic 80.18 98.83 162.27 68.11 63.52
Aquarium 122.71 98.38 146.26 162.19 83.07
Underwater 73.29 147.33 126.58 125.32 70.20
Table 2: mAP@.50:.95 ($\uparrow$) of YOLOv5s / YOLOv7 on RF7. Baseline models are trained with 200 real images only, whereas the other models are trained with 200 real + 5000 synthetic images from various methods. ODGEN leads to the biggest improvement on all 7 domain-specific datasets.
Baseline ReCo GLIGEN ControlNet GeoDiffusion ODGEN
real + synth # 200 + 0 200 + 5000 200 + 5000 200 + 5000 200 + 5000 200 + 5000
Apex Game 38.3 / 47.2 25.0 / 31.5 24.8 / 32.5 33.8 / 42.7 29.2 / 35.8 39.9 / 52.6
Robomaster 27.2 / 26.5 18.2 / 27.9 19.1 / 25.0 24.4 / 32.9 18.2 / 22.6 39.6 / 34.7
MRI Image 37.6 / 27.4 42.7 / 38.3 32.3 / 25.9 44.7 / 37.2 42.0 / 38.9 46.1 / 41.5
Cotton 16.7 / 20.5 29.3/ 37.5 28.0 / 39.0 22.6 / 35.1 30.2 / 36.0 42.0 / 43.2
Road Traffic 35.3 / 41.0 22.8 / 29.3 22.2 / 29.5 22.1 / 30.5 17.2 / 29.4 39.2 / 43.8
Aquarium 30.0 / 29.6 23.8 / 34.3 24.1 / 32.2 18.2 / 25.6 21.6 / 30.9 32.2 / 38.5
Underwater 16.7 / 19.4 13.7 / 15.8 14.9 / 18.5 15.5 / 17.8 13.8 / 17.2 19.2 / 22.0
4 Experiments
We conduct extensive experiments to demonstrate the effectiveness of the proposed ODGEN in both specific and general domains. FID [18] is used as the metric of fidelity. Mean average precisions (mAP) of YOLOv5s [23] and YOLOv7 [52] trained with synthetic data are used to evaluate the trainability, which concerns the usefulness of synthetic data for detector training. Our approach is implemented with Stable Diffusion v2.1 and compared with prior controllable generation methods based on Stable Diffusion, including ReCo [59], GLIGEN [31], ControlNet [60], GeoDiffusion [5], and MIGC [65]. Native ControlNet doesn’t support bounding box conditions. Therefore, we convert boxes to a mask $C$ sized $H\times W\times K$, where $H$ and $W$ are the image height and width, and $K$ is the class number. If one pixel $(i,j)$ is in $n$ boxes of class $k$, the $C[i,j,k]=n$. YOLO models are trained with the same recipe as Roboflow-100 [6] for 100 epochs to ensure convergence. The length of text and image lists $N$ is set to the maximum object number per image in the training set.
4.1 Specific Domains
Roboflow-100 [6] is used for the evaluation in specific domains. It consists of 100 object detection datasets of various domains. We select 7 representative datasets (RF7) covering video games, medical imaging, and underwater scenes. To simulate data scarcity, we only sample 200 images as the training set for each dataset. We fine-tune all methods with the 200 images and generate 5000 new synthetic images for evaluation. $\lambda$ in Eq. 1 and $\gamma$ in Eq. 2 are set as 1 and 25. We first fine-tune the diffusion model according to Fig. 2 (a) for 3k iterations. Then we train the object-wise conditioning modules on a V100 GPU with batch size 4 for 200 epochs, the same as the other methods to be compared.
4.1.1 Fidelity
For each dataset in RF7, we compute the FID scores of synthetic images against real images, respectively. As shown in Tab. 1, our approach outperforms all the other methods. We visualize the generation quality in Fig. 4. GLIGEN only updates gated self-attention layers inserted to Stable Diffusion, making it hard to be generalized to specific domains like MRI images and the Apex game. ReCo and GeoDiffusion integrate encoded bounding box information into text tokens, which require more data for diffusion model and text encoder fine-tuning. With only 200 images, they fail to generate objects in the given box regions. ControlNet, integrating the bounding box information into the visual condition, presents more consistent results with the annotation. However, it may still miss some objects or generate wrong categories in complex scenes. ODGEN achieves superior performance on both visual effects and consistency with annotations.
4.1.2 Trainability
To explore the effectiveness of synthetic data for detector training, we train YOLO models pre-trained on COCO with different data and compare their mAP@.50:.95 scores in Tab. 2. In the baseline setting, we only use the 200 real images for training. For the other settings, we use a combination of 200 real and 5000 synthetic images. Our approach gains improvement over the baseline and outperforms all the other methods.
4.2 General Domains
COCO-2014 [35] is used to evaluate our method in general domains. As the scenes in this dataset are almost covered by the pre-training data of Stable Diffusion, we skip the domain-specific fine-tuning stage and train the diffusion model together with the object-wise conditioning modules. We train our ODGEN on the COCO training set with batch size 32 for 60 epochs on ×$\times 8$ V100 GPUs as well as GLIGEN, ControlNet, and GeoDiffusion. Officially released checkpoints for ReCo (trained for 100 epochs) and MIGC (trained for 300 epochs) are directly employed for comparison. $\gamma$ in Eq. 2 is set as 10 for our ODGEN trained on COCO.
Figure 5: Visualized results comparison for models trained on COCO. ODGEN is better qualified for synthesizing complex scenes with multiple categories of objects and bounding box occlusions.
Table 3: FID($\downarrow$) and mAP($\uparrow$) of YOLOv5s / YOLOv7 on COCO. FID is computed with 41k synthetic images. For mAP, YOLO models are trained from scratch on 10k synthetic images and validated on 31k real images. ODGEN outperforms all the other methods in terms of both fidelity and trainability.
Metrics ReCo GLIGEN ControlNet GeoDiffusion MIGC ODGEN
FID 18.36 26.15 25.54 30.00 21.82 16.16
mAP@.50 7.60 / 11.01 6.70 / 9.42 1.43 / 1.15 5.94 / 9.21 9.54 / 16.01 18.90 / 24.40
mAP@.50:.95 3.82 / 5.29 3.56 / 4.60 0.52 / 0.38 2.37 / 4.44 4.67 / 8.65 9.70 / 14.20
4.2.1 Fidelity
We use the annotations of the COCO validation set (41k) as conditions to generate 41k images. The FID scores in Tab. 3 are computed against real images in the COCO validation set. ODGEN achieves better FID results than the other methods. We provide a typical visualized example in Fig. 5. Given overlapping bounding boxes of multiple categories, ODGEN synthesizes all the objects of correct categories and accurate positions, outperforming the other methods. Substantial visualized samples are added in Appendix F.
4.2.2 Trainability
We use 10k annotations randomly sampled from the COCO validation set to generate a synthetic dataset of 10k images. The YOLO models are trained on this synthetic dataset from scratch and evaluated on the other 31k images in the COCO validation set. As shown in Tab. 3, ODGEN achieves significant improvement over the other methods in terms of mAP.
Table 4: Using the image list (IL) and the text list (TL) benefits FID and mAP (YOLOv5s / YOLOv7). They are especially helpful for the Road Traffic dataset which has more categories and occlusions.
Datasets IL TL FID($\downarrow$) mAP@.50:.95 ($\uparrow$)
MRI Image $\times$ $\times$ 98.29 44.6 / 39.9
$\times$ $\checkmark$ 95.67 44.9 / 41.4
$\checkmark$ $\times$ 99.16 46.2 / 41.0
$\checkmark$ $\checkmark$ 93.82 46.1 / 41.5
Road Traffic $\times$ $\times$ 67.40 25.5 / 35.4
$\times$ $\checkmark$ 66.48 26.5 / 37.0
$\checkmark$ $\times$ 65.80 32.3 / 39.1
$\checkmark$ $\checkmark$ 63.52 39.2 / 43.8
Table 5: Proper $\gamma$ values in Eq. 2 benefit both FID and mAP (YOLOv5s / YOLOv7) of synthetic images, while overwhelming values lead to degeneration.
Datasets $\gamma$ FID ($\downarrow$) mAP@.50:.95 ($\uparrow$)
Aqua- rium 0 83.94 30.8 / 34.5
25 83.07 32.2 / 38.5
50 87.00 32.1 / 38.0
100 90.13 29.9 / 35.7
Road Traffic 0 66.65 36.9 / 39.5
25 63.52 39.2 / 43.8
50 66.46 36.6 / 38.9
100 72.18 32.9 / 37.1
Table 6: mAP (YOLOv5s / YOLOv7) of the corrupted label filtering for ODGEN.
Datasets Label filtering mAP@.50:.95 ($\uparrow$)
Cotton $\checkmark$ 42.0 / 43.2
$\times$ 40.5 / 42.1
Robo- master $\checkmark$ 39.6 / 34.7
$\times$ 39.0 / 33.3
Under- water $\checkmark$ 19.2 / 22.0
$\times$ 18.9 / 21.6
4.3 Ablation Study
Image list and text list in object-wise conditioning modules. ODGEN enables object-wise conditioning with an image list and a text list as shown in Fig. 2. We test the performance of using global text prompts in the placement of text lists or pasting all the foreground patches on the same empty canvas in the placement of image lists for the ablation study. Experiments are performed on the MRI and Road Traffic datasets in RF7. The MRI dataset only has two categories and almost no overlapping objects, whereas the road traffic dataset has 7 categories and many occluded objects. From the results in Tab. 5, it can be found that using image lists or text lists brings moderate improvement on MRI images and significant improvement on the more complex Road Traffic dataset. Visualization in Fig. 6 shows that the image list improves the fidelity under occlusion, and the text list mitigates the mutual interference of multiple objects.
Foreground region enhancement. We introduce a re-weighting term controlled by $\gamma$ for foreground objects in Eq. 2.  Tab. 5 shows that appropriate $\gamma$ values can improve both the fidelity and the trainability since the foreground details are better generated. However, increasing the value will undermine the background quality, as visualized in Fig. 6.
Corrupted label filtering. Results in Tab. 6 show that the corrupted label filtering step helps improve the mAP@.50:95 of ODGEN by around 1-2%. The results of ODGEN without the corrupted label filtering still significantly outperform the baseline setting in Tab. 2.
Figure 6: Visualized ablations. Left: using neither image nor text lists, a traffic light is generated in the position of a bus, and a car is generated in the position of a traffic light; Middle: not using image list, two occluded motorcycles are merged as one; Right: not using text list, a motorcycle is generated in the position of a vehicle. Using both image and text lists generates correct results.
Figure 7: With increasing $\gamma$ value, our approach produces higher-quality foreground objects. However, overwhelming $\gamma$ leads to blurred background and degenerated fidelity.
5 Conclusion
This paper introduces ODGEN, a novel approach aimed at generating domain-specific object detection datasets to enhance the performance of detection models. We propose to fine-tune the diffusion model on both cropped foreground objects and entire images. We design a mechanism to control the foreground objects with object-wise synthesized visual prompts and textual descriptions. Our work significantly enhances the performance of diffusion models in synthesizing complex scenes characterized by multiple categories of objects and bounding box occlusions. Extensive experiments demonstrate the superiority of ODGEN in terms of fidelity and trainability across both specific and general domains. We believe this work has taken an essential step toward more robust image synthesis and highlights the potential of benefiting the object detection task with synthetic data.
Appendix A Limitations
Despite the significant improvement achieved by this work, we find it remains challenging to achieve comparable training performance as real data with merely synthetic data. We consider the limitation to be mainly caused by two reasons. Firstly, the generation fidelity is still limited by modern diffusion models. For instance, VAEs in Stable Diffusion are not good enough to deal with complex text in images, constraining the performance on video game datasets. ODGEN is promising to achieve further improvements with more powerful diffusion models in the future. Secondly, this paper mainly focuses on designing generative models with given conditions. However, the other parts in the dataset synthesis pipeline are not fully optimized, which are interesting to be explored in future work.
Appendix B Broader Impact
We introduce a novel approach for object detection datasets synthesis, achieving improvement in fidelity and trainability for both specific and general domains. Our approach is more prone to biases caused by training data than typical AI generative models since it shows robustness with limited data of specific domains. Therefore, we recommend practitioners to apply abundant caution when dealing with sensitive applications to avoid problems of races, skin tones, or gender identities.
Appendix C Supplemental Ablations
Detectors training w/o real data. In the main paper, we train the detectors with 200 real images and 5000 synthetic images for RF7 benchmarks. In this ablation study, we conduct experiments with the 5000 synthetic images only, to explore the impact of the absence of real data for training detectors. As shown in Tab. 7, for easy datasets with fewer objects, like MRI image and Cotton, training with 5000 synthetic data only achieves better results than training with 200 real samples only. Because the generation quality is good enough in these cases, increasing the dataset size can improve the model performance. Nevertheless, for complex scenes like Robomaster and Aquarium, training with synthetic data only fails to achieve better results. It is caused by the limited generation quality of modern diffusion models. For instance, the VAE in Stable Diffusion is not good at synthesizing texts in images, which are common in the Robomaster dataset. Besides, FID results in Tab. 1 indicate that there still exist gaps between the distributions of real and synthetic images. The performance of our approach is promising to be further improved in the future with more powerful generative diffusion models.
Table 7: Ablations of using real data only and using synthetic data only during detector training. (Metric: YOLOv5s / YOLOv7)
Metrics mAP@.50 ($\uparrow$) mAP@.50:.95 ($\uparrow$)
real + synth # 200 + 0 0 + 5000 200 + 5000 200 + 0 0 + 5000 200 + 5000
Robomaster 50.3 / 48.8 35.9 /37.2 67.4 / 65.4 27.2 / 26.5 18.7 / 21.0 39.6 / 34.7
MRI Image 55.1 / 44.7 59.8 / 55.2 68.1 / 61.3 37.6 / 27.4 41.3 / 35.3 46.1 / 41.5
Cotton 29.4 / 32.0 43.1 / 43.2 63.9 / 55.0 16.7 / 20.5 29.5 / 28.1 42.0 / 43.2
Aquarium 53.8 / 53.0 45.2 / 41.1 62.6 / 66.5 30.0 / 29.6 23.8 / 20.9 32.2 / 38.5
Category isolation for datasets synthesis. As illustrated in the main paper, modern diffusion models have the limitation of "concept bleeding" when multiple categories are involved in the generation of one image. An alternative method could be only generating objects of the same category in one image. We select 3 multi-class datasets and compare this approach against the proposed method in ODGEN. As shown in Tab. 8, ODGEN achieves better results on most benchmarks than generating samples with isolated categories. This is an expected result since generating objects of various categories in one image should be better aligned with the test set.
Table 8: Ablations of category isolation during datasets synthesis (Metric: mAP@.50 of YOLOv5s / YOLOv7, real + synth images # 200 + 5000). It hinders the performance in most cases.
Category Isolation True False
Road Traffic 63.4 / 70.3 66.8 / 70.4
Aquarium 57.2 / 63.3 62.6 / 66.5
Underwater 41.1 / 42.0 40.0 / 44.8
Appendix D Implementation Details
D.1 Global Text Prompt
All the methods included for comparison in our experiments share the same textual descriptions of the foreground objects and entire scenes in the global text prompts for a fair comparison. ReCo [59] and GeoDiffusion [5] further add the location information of foreground objects to their global text prompts following their approaches. For general domains, we concatenate the object class names into a sentence as the global text prompt. For specific domains, we concatenate the object class names and the scene name into a sentence as the global text prompt.
D.2 ODGEN (ours)
We fine-tune the parameters in the UNet of Stable Diffusion with the default training setting used by ControlNet [60], including the learning rates and the optimizer.
Offline foreground objects synthesis. To improve the training efficiency, we synthesize 500 foreground object images for each category offline and build a synthetic foreground pool. During the training of object-wise conditioning modules, foreground images are randomly sampled from the pool to build the image list.
Image encoder. ControlNet uses a 4-layer convolutional encoder to encode input conditions sized $512$ to $64\times 64$ latents. The input channels are usually set as 1 or 3 for input images. Our ODGEN follows similar methods to convert the input conditions. The key difference is that we use image lists of synthesized objects as input conditions. The image lists are padded to $N$ and concatenated in the channel dimension for encoding. In this case, we expand the input channels of the encoder to $3N$ and update the channels of the following layers to comparable numbers. We provide the channels of each layer in Tab. 9. The convolutional kernel size is set as 3 for all layers.
Table 9: Detailed input and output channels of each convolutional layer in the image encoder used in ODGEN. As the maximum object number per image for different datasets varies, we select different $N$ and thus different channel numbers.
Datasets $N$ Layer 1 Layer 2 Layer 3 Layer 4
Input Output Input Output Input Output Input Output
Apex Game 6 18 32 32 96 96 128 128 256
Robomaster 17 51 64 64 96 96 128 128 256
MRI Image 2 6 16 16 32 32 96 96 256
Cotton 27 81 96 96 128 128 192 192 256
Road Traffic 19 57 64 64 96 96 128 128 256
Aquarium 56 168 168 168 192 192 224 224 256
Underwater 79 237 237 237 256 256 256 256 256
COCO 93 297 297 297 256 256 256 256 256
[…]
Foreground/Background discriminator. We train discriminators by fine-tuning ImageNet pre-trained ResNet50 on foreground and background patches randomly cropped from training sets. They are split for training, validating, and testing by 70%, 10%, and 20%. The model is a binary classification model that only predicts whether a patch contains any object. The accuracy on test datasets is over 99% on RF7. Therefore, we can confidently use it to filter the pseudo labels.
D.3 Other Methods
ReCo [59] & MIGC [65] & GeoDiffusion [5] are layout-to-image generation methods based on Stable Diffusion. For the comparison on RF7, we use their official codes to fine-tune the diffusion models on domain-specific datasets and generate synthetic datasets for experiments. For the comparison on COCO, we directly use their released pre-trained weights trained on COCO. MIGC only releases the pre-trained weight but not the code, so we only compare it with our method on the COCO benchmark. ReCo and GeoDiffusion fine-tune both the text encoder and the UNet of Stable Diffusion. MIGC only fine-tunes the UNet.
GLIGEN [31] is an open-set grounded text-to-image generation methods supporting multiple kinds of conditions. We use GLIGEN for layout-to-image generation with bounding boxes as conditions. GLIGEN is designed to fit target distributions by only fine-tuning the inserted gated self-attention layers in pre-trained diffusion models. Therefore, we fix the other layers in the diffusion model during fine-tuning on RF7 or COCO.
ControlNet [60] is initially designed to control diffusion models by fixing pre-trained models, reusing the diffusion model encoder and fine-tuning it to provide control. In our experiments, to make fair comparisons, we also fine-tune the diffusion model part to fit it to specific domains that can be different from the pre-training data of Stable Diffusion. The diffusion model remains trainable for experiments on COCO as well. The native ControlNet is originally designed as an image-to-image generation method. As illustrated in Sec. 4, we provide a simple way to convert boxes to masks. Given a dataset containing $K$ categories, an image sized $H\times W$, and $B$ objects with their bounding boxes $bbox$ and classes $cls$, we convert them to a mask $M$ in the following process:
Algorithm 1 Convert bounding boxes to a mask for ControlNet
  $M$ = zeros($H$, $W$, $K$)
  for $i=1,\dots,B$ do
     $left,top,width,height=bbox[i]$
     $k=cls[i]$
     $M[top:top+height,left:left+width,k]+=1$
  end for
Table 10: The number of images in the 7 subsets of Roboflow-100 used to compose the RF7 datasets.
Datasets Train Validation Test
Apex Game 2583 415 691
Robomaster 1945 278 556
MRI Image 253 39 79
Cotton 367 20 19
Road Traffic 494 133 187
Aquarium 448 63 127
Underwater 5320 760 1520
Appendix E Datasets Details
RF7 datasets. We employ the RF7 datasets from Roboflow-100 [6] to evaluate the performance of our ODGEN on specific domains. Here, we add more details of the employed datasets, including Apex Game, Robomaster, MRI Image, Cotton, Road Traffic, Aquarium, and Underwater. We provide the number of images for each dataset, including the train, validation, and test sets in Tab. 10. We use the first 200 samples in the image list annotation provided by Roboflow-100 for each dataset to build the training set for our experiments. The full validation set is used as the standard to choose the best checkpoint from YOLO models trained for 100 epochs. The full test set is used to evaluate the chosen checkpoint and provide mAP results shown in this paper. Visualized examples with annotations are shown in Fig. 8. The object categories of each dataset are provided in Tab. 11. Some images in the Apex Game and Robomaster datasets contain complex and tiny text that are hard to generate by current Stable Diffusion.
Figure 8: Visualized examples with object detection annotations of RF7 datasets.
Table 11: Class names of the RF7 datasets.
Datasets Class Names
Apex Game Avatar, Object
Robomaster Armor, Base, Car, Rune, Rune-blue, Rune-gray, Rune-grey, Rune-red, Watcher
MRI Image Negative, Positive
Cotton G-arboreum, G-barbadense, G-herbaceum, G-hirsitum
Road Traffic Bicycles, Buses, Crosswalks, Fire hydrants, Motorcycles, Traffic lights, Vehicles
Aquarium Fish, Jellyfish, Penguin, Puffin, Shark, Starfish, Stingray
Underwater Echinus, Holothurian, Scallop, Starfish, Waterweeds
Table 12: mAP@.50 ($\uparrow$) of YOLOv5s / YOLOv7 on RF7. Baseline models are trained with 200 real images only, whereas the other models are trained with 200 real + 5000 synthetic images from various methods. ODGEN leads to the biggest improvement on all 7 domain-specific datasets.
Baseline ReCo GLIGEN ControlNet GeoDiffusion ODGEN
real + synth # 200 + 0 200 + 5000 200 + 5000 200 + 5000 200 + 5000 200 + 5000
Apex Game 64.4 / 78.5 53.3 / 60.0 53.3 / 57.0 59.3 / 70.3 55.8 / 62.1 69.1 / 82.6
Robomaster 50.3 / 48.8 41.6 / 54.4 38.7 / 50.2 44.4 / 57.0 41.5 / 45.7 67.4 / 65.4
MRI Image 55.1 / 44.7 59.7 / 59.6 52.1 / 42.0 64.7 / 55.4 61.6 / 56.1 68.1 / 61.3
Cotton 29.4 / 32.0 42.7 / 53.4 41.8 / 55.0 38.3 / 47.1 49.6 / 49.4 63.9 / 55.0
Traffic 61.1 / 68.8 44.6 / 52.4 45.0 / 54.5 43.5 / 56.2 38.2 / 54.6 66.8 / 70.4
Aquarium 53.8 / 53.0 45.8 / 59.0 49.7 / 51.2 35.5 / 45.4 45.1 / 54.8 62.6 / 66.5
Underwater 35.7 / 39.2 30.9 / 34.5 33.6 / 38.7 34.3 / 37.7 31.8 / 37.6 40.0 / 44.8
Figure 9: Additional visualized samples produced by our approach in various specific domains. Annotations of bounding boxes are marked on samples and global text prompts are provided below corresponding samples. Our approach can be generalized to various domains and is compatible with complex scenes with multiple objects, multiple categories, and bounding box occlusions.
Figure 10: Visualized comparison results on the COCO-2014 dataset. The annotations used as conditions to synthesize images are shown in the first column.
Figure 11: Visualized comparison results on the COCO-2014 dataset. The annotations used as conditions to synthesize images are shown in the first column.
Figure 12: Visualized comparison results on the COCO-2014 dataset. The annotations used as conditions to synthesize images are shown in the first column.
Figure 13: Visualized results of ODGEN on the COCO-2014 dataset. In every grid, the annotation used as conditions to synthesize images are shown in the top-left corner.
Figure 14: Visualized results of our ODGEN on the COCO-2014 dataset generated from the same conditions. The annotation is placed on the first column.
Appendix F Supplemental Experiments
Other related works. InstaGEN [11] is mainly designed for the open-vocabulary object detection task. It generates images randomly and annotates them with a grounding head trained with another pre-trained detector. InstanceDiffusion [53] needs detailed segmentation masks for training and inference. They are not included for direct comparison in our paper.
mAP@.50 on RF7.  Tab. 12 provides quantitative results for trainability in terms of mAP@.50 as supplements to Tab. 2. Our approach achieves improvement over training on real data only and outperforms other methods.
Visualization. We provide extensive visualized examples of ODGEN to demonstrate its effectiveness. In Fig. 9, we show generated samples of the specific domains in RF7. It can be seen that ODGEN is robust to complex scenes composed of multiple categories, dense and overlapping objects. Besides, ODGEN shows strong generalization capability to various domains.
In Fig. 10Fig. 11, and Fig. 12, we provide supplemental visualized comparisons between ODGEN and the other methods on the COCO-2014 dataset. ODGEN achieves the best image-annotation consistency and maintains high generation fidelity. More generated samples of ODGEN are shown in Fig. 13. In addition, we also show different examples produced by ODGEN on the same conditions in Fig. 14 to show its capability of generating diverse results.